jakfishman:

Two lines, E and F, are represented by the equations given below. Line E: 5x + 5y = 40 Line F: x + y = 8 Which statement is true about the solution to the set of equations? It is (40, 8). It is (8, 40). There is no solution. There are infinitely many solutions.

jakfishman:

@supie sorry for @ you but do you know?

jakfishman:

@jhonyy9 sorry for @ you but do you know?

jhonyy9:

ok like a first step divide both sides by 5 in case of the first equation - so what will get ?

jakfishman:

x=y?

jhonyy9:

no

jhonyy9:

5x +5y = 40 5x +5y = 8*5 5(x +y) = 8*5 divide both sides by 5 x+y = 8 so what mean this ?

jakfishman:

1.6 and 0?

StrappedQuez:

4 bc 5x4 is 20 so add those together get 40 and 4 + 4 = 8 so x=4 y=4

jhonyy9:

no the first equation is the same like the second - so what mean this for the system ?

jakfishman:

I'm lost

jhonyy9:

the last one choice

jhonyy9:

when in a set of two equations are two the same equations so the set are infinitely many solutions

jhonyy9:

@Laylalyssa pls